OpenStudy (anonymous):

solve the system of linear equations 4x-3y=17 5x+4y=60

OpenStudy (anonymous):

Same as the previous one i helped you with but you need to multiply both lines by 2 different numbers the first line being 4 and 2nd line being -3. so you get 16x-12y=68 -15x+12y=-180 and you get x=-112 then plug -112 back into one of the 2 (i will do the first line) and you get 16(-112)-12y=68 -1792-12y=68 -12y=1860 y=-155 then you can check by plugging both values in: 16(-112)-12(-155)=68 -15(-112)+12(-155)=-180